OpenStudy (anonymous):

What is the value of the coefficient "b" when the quadratic equation y = (4x + 1)(2x - 3) is written in standard form?

OpenStudy (anonymous):

well u gotta foil it and u will land up with an equation: y=ax^2+bx+c y=8x^2-12x+2x-3 y=8x^2-10x-3 b=-10 If i did my arithmetic correctly